Vicky Steeves

� (212) 992 6269 � vicky.steeves@nyu.edu � vickysteeves.com � VickySteeves � VickySteeves � VickySteeves

Objective

I work to support researchers in creating well-managed, high quality, and reproducible research through integrating reproducible practices into the research workflow, advocating openness in all facets of scholarship, and building/contributing to open infrastructure.

Education

Simmons College  Boston, MA
Master of Library & Information Science  2013 – 2014
Simmons College  Boston, MA
Bachelor of Science Computer Science & Information Technology  2010 – 2013

Graduate and Undergraduate Research Experience. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Supervisor: Kathy Wisser  Boston, MA
Small World Project  Spring 2014

Analyzed research done with social media networks on archival material and visualized the analysis using Gephi.

Supervisor: Nanette Veilleux  Boston, MA
Computational Linguistic Approach to Inflection in Human Speech & Difference 2012 – 2013

Recorded speech exemplars and used a specialized programming language (Praat) to write a program to determine the tonal center of gravity of each exemplar and make sure that it matches the intended intonational contour.

University College Cork  Cork, Ireland
Celtic Studies  July 2012

Completed coursework in Irish language, literature, and history.

Awards and Honors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Simmons College

Dean’s Fellow Award  2013 – 2014

Given to students that demonstrate academic excellence.

Simmons College Computer Science & Information Technology Department

The Computer Science Award  2013

Given for excellence in Computer Science.

Simmons College

Collaborative Research Experience for Undergraduates  2012 – 2013

A research grant given to encourage women in STEM fields to complete research studies in their fields.

Simmons College

Bowker Award for Travel  2012

Given to undergraduates who demonstrate academic excellence, completing their studies abroad.

Simmons College

Simmons College 3+1 Undergraduate and Graduate Program  2010 – 2014

First participant.

Professional Experience

Simmons College  Boston, MA
Adjunct Professor  2017 –

I teach database management online for the School of Information and Library Science. Fall 2017 & Spring 2018.

New York University  New York, NY
Librarian for Research Data Management & Reproducibility  2015 –

A dual appointment between NYU’s Division of Libraries and Center for Data Science. In this role, I work to support students, faculty, staff, and researchers in creating well-managed, high quality, and reproducible research through facilitating use of tools such as ReproZip, teaching and instruction, and group and individual consultations.

Metropolitan New York Library Council  New York, NY
Interim Program Coordinator  2015 – 2016

I handled the day-to-day operations of METRO’s National Digital Stewardship Residency in New York program. I contributed to project planning, communications, documentation, evaluations, outreach, as well as maintained the program’s web presence. I planned, organized, and ran NDSR-affiliated events, meetings, and workshops.

American Museum of Natural History  New York, NY
National Digital Stewardship Resident  2014 – 2015

My project at the AMNH consisted of performing an environmental scan of the Science divisions to better understand their data storage, curation, and preservation needs. After data collection, I identified existing practices and policies for integrated data storage, access, and management. After data analysis, I recommended strategies to digitally preserve the scientific research at the AMNH.

Sasaki Associates  Waterville, MA
Archives Intern  Spring 2014

I processed historical architectural material and wrote the accompanying finding aid and created records for each, then catalogued the collections in Koha ILS

Simmons College  Boston, MA
Dean’s Fellow for Technology  2013 – 2014

I managed social media for the undergraduate science departments to generate interest in STEM at Simmons through outreach to alumnae, current students, and prospective students, working on content creation with faculty.

Other Professional Work

The LIS Scholarship Archive: Co-director of LISSA, a parntership with the Center for Open Science & the LIS community to build a place for LIS scholars to share their work. SPARC: Contributor to the SPARC International Survey on Research Data Management Readiness (with Nick Wolf and Scott Collard). Radian: Editor. The Idealis: Editor. CLIR: Participant, 2016 CLIR eResearch Network Research Data Management cohort. Preservation and Archiving Special Interest Group: Chaired the committee responsible for the reproducibility track of the Fall 2016 conference. Center for Open Science: Open Science Framework Ambassador. Participant in the 2016 SHARE Curation Associates program. DLF: Participant in the 2016 eDLF Data Management Cohort.

Data And Software Preservation for Open Science: External organizer for the workshop on containerizing science. Library Information and Technology Association: Taught a one-month course on building library services around openness and reproducibility. Preservation and Quality Tools: Facilitated a hands-on session at this workshop on ReproZip. Data And Software Preservation for Open Science: Led multiple hands-on sessions on ReproZip. Presented to the DASPOS board of directors on NYU’s open source reproducibility initiatives. OpenCon: Guest lectured on reproducibility for the OpenCon early career librarian’s call. Metropolitan NY Library Council (METRO): Presented a webinar teaching RDM principles to library and information professionals. Pratt School of Library and Information Science: Guest lectured in Meg Smith’s Data Librarianship class on my career path, RDM, and reproducibility. 2016 & 2017 Palmer School of Library & Information Science: Guest lectured in Don Mennerich’s Digital Archives class for the Public History program on RDM. 2016 & 2017 Queen’s College Graduate School of Library & Information Science: Guest lectured in Robert DeCandido’s digital stewardship class at Queens College on data management, NDSR, and my career path.

2015 – 2017 Works

Publications. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Rémi Rampin, Fernando Chirigati, Vicky Steeves, and Juliana Freire. “ReproServer: Making Reproducibility Easier and Less Intensive.” arXiv:1808.01406 [cs], aug 2018. Available at

https://arxiv.org/abs/1808.01406.

Vicky Steeves and Kelly Johnson. “Research data management among life sciences faculty.” The LIS Scholarship Archive, 2018. doi:10.17605/OSF.IO/Q36UV.

Vicky Steeves, Rémi Rampin, and Fernando Chirigati. “Using ReproZip for reproducibility and library services.” IASSIST Quarterly, 42(1):14, dec 2017. doi:10.29173/iq18. Avail-

able at https://iassistquarterly.com/index.php/iassist/article/view/18. Preprint at
https://osf.io/preprints/lissa/5tm8d/.

Vicky Steeves.  “Reproducibility Librarianship.”  Collaborative Li-
brarianship,  9(2),  jul 2017.  ISSN 1943-7528.  Available at

http://digitalcommons.du.edu/collaborativelibrarianship/vol9/iss2/4.

Rémi Rampin, Fernando Chirigati, Vicky Steeves, Dennis Shasha, and Juliana Freire. “ReproZip: The Reproducibility Packer.” The Journal of Open Source Software, 1(8), dec 2016. doi:10.21105/

joss.00107. Available at http://joss.theoj.org/papers/10.21105/joss.00107.

Karl-Rainer Blumenthal, Peggy Griesinger, Julia Kim, Shira Peltzman, and Vicky Steeves. “What makes a digital steward: A competency profile based on the National Digital Stewardship Residencies.” In “iPres: 13th International Conference on Digital Preservation,” 2016. doi:

10.17605/OSF.IO/TNMRA. Available at https://osf.io/preprints/lissa/tnmra/.

Andreas Wolke, Martin Bichler, Fernando Chirigati, and Victoria Steeves. “Reproducible experiments on dynamic resource allocation in cloud data centers.” Information Systems, 2016. ISSN 0306-4379. doi:10.1016/j.is.2015.12.004. Postprint available at https://arxiv.org/abs/1703.00042.

Presentations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Vicky Steeves, Rémi Rampin, Fernando Chirigati, and Juliana Freire. “Reproserver: Making reproducibility easier and less intrusive.” In “MSDSE Summit 2018,” 2018. Available at

https://osf.io/6hvpt.

Vicky Steeves. “Best practices for reproducible research.” In “Harvard-Smithsonian center for Astrophysics,” 2018. Available at https://osf.io/6hvpt.

Vicky Steeves. “Reproducing & archiving scholarship with reprozip.” In “PresQT Workshop Two,” 2018. doi:10.17605/OSF.IO/S6FW9. Available at https://osf.io/tavwn/.

Vicky Steeves. “Research with interviews.” In “NDSR-Art Immersion Week,” 2018. Available at

https://vickysteeves.gitlab.io/2018-ndsr-art/.

Vicky Steeves. “Reproducibility short course.” In “Building Integrity through Reproducibility,” 2018.

Available at https://vickysteeves.gitlab.io/2018-uutah-repro/.

Vicky Steeves, Daniel Nüst, Markus Konkel, Rémi Rampin, and Edzer Pebesma. “Writing reprocible geoscience papers using r markdown, docker, and gitlab.” In “European Geosciences Union,” 2018.

Available at https://vickysteeves.gitlab.io/repro-papers/.

Vicky Steeves. “Data management and reproducibility.” In “L’Unité Régionale de Formation Ãă l’Information Scientifique et Technique,” 2017. Available at https://osf.io/6fyj5/.

Vicky Steeves and Rémi Rampin. “Creating reproducible experiments with reprozip.” In “SciPy

2017,” 2017. Available at https://vickysteeves.gitlab.io/2017-SciPy/#/.

Vicky Steeves, Cynthia Hudson-Vitale, and Matt Spitzer. “Open, public goods infrastructure for research management & discovery.” In “IASSIST 2017,” 2017. doi:10.17605/OSF.IO/SY2ZF.

Available at https://osf.io/sy2zf/.

Vicky Steeves, Rémi Rampin, and Fernando Chirigati. “Reproducing and pre-

serving research with reprozip.”  In “IASSIST 2017,” 2017.  Available at

https://vickysteeves.gitlab.io/2017-IASSIST-ReproZip/.

Vicky Steeves and Rémi Rampin. “Reprozip for reproducible research.” In “PresQT Workshop One,” 2017. doi:10.17605/OSF.IO/S6FW9. Available at https://osf.io/s6fw9/.

Vicky Steeves, Rémi Rampin, and Daniel Nüst. “Reproducible computational research in the publication cycle.” In “European Geosciences Union General Assembly 2017,” 2017. doi:10.17605/ OSF.IO/UMY6G. Available at https://osf.io/umy6g/.

Vicky Steeves. “Open science: Understanding modern research practices.” In “Association of College and Research Libraries,” 2017. doi:10.17605/osf.io/cyr3n. Available at

https://osf.io/preprints/lissa/cyr3n/.

Vicky Steeves and Nicholas Wolf. “Using Openness as Foundation for Research Data Management Services.” In “Library Information and Technology Association Conference,” 2016. Available at

http://forum.lita.org/sessions/using-openness-as-foundation-for-research-data-management-services/.

Karl-Rainer Blumenthal, Peggy Griesinger, Julia Kim, Shira Peltzman, and Vicky Steeves. “What makes a digital steward: A competency profile based on the National Digital Stewardship Residencies.” In “iPres: 13th International Conference on Digital Preservation,” 2016. doi: 10.17605/OSF.IO/XFC26. Available at https://osf.io/xfc26/.

Vicky Steeves and Rémi Rampin. “ReproZip: Reproducibility with Ease.” In “DASPOS: Container Strategies for Data & Software Preservation that Promote Open Science,” 2016. doi:10.17605/ OSF.IO/AQZTF. Available at https://osf.io/aqztf/.

Vicky Steeves, Kevin B. Read, and Andrew S. Gordon. “Collaborating to Create a Culture of Data Stewardship.” In “Research Data Access and Preservation Summit,” 2016. Available at

http://hdl.handle.net/2451/38185.

Vicky Steeves, Kevin B. Read, and Andrew S. Gordon. “Bridging the Gap: Improving Data Services through Cross-Campus Collaboration.” In “University of Massachusetts and New England Area Librarian e-Science Symposium,” 2016. Available at

http://escholarship.umassmed.edu/escience_symposium/2016/posters/2/.

Vicky Steeves, Jennifer Muilenburg, and Erik Mitchell. “Organizational implications of data science environments in education, research, and research management in libraries.” In “Coalition for Networked Information Fall Meeting,” 2015. doi:http://doi.org/10.17605/OSF.IO/RHGQ3.

Available at https://osf.io/rhgq3/.

Code & Media. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Rémi Rampin, Vicky Steeves, and Sarah DeMott. “Taguette: Floss qualitative research tool.”, 2018.

Available at https://gitlab.com/remram44/taguette.

Vicky Steeves and David Millman. “Investigating & archiving git.”, 2018. Sloan grant to investigate options for archiving scholarship in git and on git-hosting platforms.

Rémi Rampin, Vicky Steeves, Fernando Chirigati, Meredith Broussard, and Katherine Boss. “Saving data journalism.”, 2018. IMLS grant to archive dynamic websites. Available at

https://savingjournalism.reprozip.org.

Vicky Steeves. “Database: Women leaders in openness.”, 2017. doi:10.17605/osf.io/5s2g6. Avail-

able at https://osf.io/preprints/lissa/5s2g6/.

Vicky Steeves and Rémi Rampin. “Reprocrawl.”, 2018. A search engine for materials relating to

reproducibility. Source: https://gitlab.com/vickysteeves/reprocrawl.

Vicky Steeves and Rémi Rampin. “Reprozip examples website.”, 2017. Available at

https://github.com/ViDA-NYU/reprozip-examples.

Vicky Steeves and Nicholas Wolf. “Research data management badge.”, 2016. Available at

https://osf.io/85bnv.

Vicky Steeves and Rémi Rampin. “Research applications & python scripts.”, 2016. Available at

https://github.com/NYU-DataServices/ManagingData-Python.

Vicky Steeves and Rémi Rampin. “Reproducibility news feed.”, 2015. Available at

https://github.com/ViDA-NYU/reproducibility-news.

Vicky Steeves, Nicholas Wolf, and Stephen Balogh. “Data dispatch.”, Sep 2016. Available at

https://osf.io/4tdms/.

Vicky Steeves and Rémi Rampin. “Reproducibile science website.”, 2017. Available at

https://github.com/ViDA-NYU/reproducible-science.

Vicky Steeves and Rémi Rampin. “Reproducibile science nyu website.”, 2017. Available at

https://github.com/ViDA-NYU/reproducible-science-nyu.

Vicky Steeves, Rémi Rampin, and Fernando Chirigati. “Packing and unpacking experiments with

reprozip.”, Nov 2015. Available at https://www.youtube.com/watch?v=-zLPuwCHXo0.

Grants Funded. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Vicky Steeves and David Millman. “Investigating & archiving git.”, 2018. Sloan grant to investigate options for archiving scholarship in git and on git-hosting platforms.

Rémi Rampin, Vicky Steeves, Fernando Chirigati, Meredith Broussard, and Katherine Boss. “Saving data journalism.”, 2018. IMLS grant to archive dynamic websites. Available at

https://savingjournalism.reprozip.org.

Technical skills

Programming Languages: Python, R, Java, C, C++, Objective C, Perl, PHP, SQL, HTML, CSS, JavaScript, jQuery, AJAX, and Praat Operating Systems: Ubuntu Linux, Microsoft Windows, Mac OSX Library Standards: XML,JSON, XSLT, DDI, EAD, DACS, PREMIS, MARC, Dublin Core

Employment History

New York University  New York, NY
Librarian for Research Data Management & Reproducibility  2015 –
Metropolitan New York Library Council  New York, NY
Interim Program Coordinator  2015 – 2016
American Museum of Natural History  New York, NY
National Digital Stewardship Resident  2014 – 2015
Simmons College  Boston, MA
Dean’s Fellow for Technology  2013 – 2014
Simmons College  Boston, MA
Technical Resource Assistant  2013 – 2014

Provide technical instruction to students, staff, and faculty at the Graduate School of Library and Information Science (GSLIS).Troubleshoot hard/software issues within GSLIS for students, staff, and faculty.

Simmons College Computer Science Department  Boston, MA
Teacher’s Assistant  2011 – 2013

Assit in creating classwork, homework, and quizzes for students to complete. Grade students’ work and report these grades to the instructor.

Simmons College Computer Science Department  Boston, MA
Lab Monitor  2011 – 2013

Assist students with troubleshooting soft/hardware issues. Repair and maintain computers in the computer science laboratory.

Research Interests

Professional Interests: Data management, reproducibility, digital preservation, digital archiving, systems analysis, database management, web development, software engineering and development, computational linguistics. Personal Interests: Animal care, outdoorswomanship, music, and creative writing.